Bryan is 25 years older than James and in 5 years Bryan will be twice as old as James, how old are they now?
Answer by ikleyn(52786)

.
Bryan is 25 years older than James and in 5 years Bryan will be twice as old as James, how old are they now?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Solution 1. System of two equations in two unknowns

B - J = 25,              (1)
B+5 = 2*(J+5).           (2)

Express B = J + 25 from (1) and substitute it into (2). You will get

(J+25) + 5 = 2*(J+5),    (3)  

or

J + 25 + 5 = 2J + 10  --->  25 + 5 -10 = 2J - J  --->  J = 20.

Answer. James is 20 years now. Brian is 20 + 25 = 45 years old now.

Check it yourself by calculating and comparing the ages in 5 years.


Solution 2. One equation in 1 unknown.

Let J be the James' age now.
Then Brian's age now is J + 25.

According to the condition,

(J+25) + 5 = 2*(J+5).    (1)

This equation is identical to the equation (3) of the Solution 1 above.

So, the solution is the same: J = 20. James is 20 years old now.

Then Brian is 20 + 25 = 45 years old now.

You have the same answer.